OpenStudy (anonymous):

√192 simplified. Leave answer in simple radical form.

myininaya (myininaya):

no i made a mistake

OpenStudy (anonymous):

ho ho

OpenStudy (anonymous):

I think I am suppost to leave is as √4*6*8

OpenStudy (anonymous):

\[192=2^6\times 3\]

myininaya (myininaya):

\[\sqrt{192}=\sqrt{64}\sqrt{3}=8\sqrt{3}\]

OpenStudy (anonymous):

Satellite i have to leave it in radical form

OpenStudy (anonymous):

so \[\sqrt{192}=2^3\sqrt{3}=8\sqrt{3}\]

OpenStudy (anonymous):

yes it is in simplest radical form

myininaya (myininaya):

that was just him thinking above

OpenStudy (anonymous):

could you smell the smoke from where you are?

OpenStudy (anonymous):

oh ok.

myininaya (myininaya):

he wants you to adopt the way he thinks

OpenStudy (anonymous):

thanks!